OpenStudy (anonymous):

7 * 3^3x-2 = 21. Can anyone please answer this question with their work.. Thanks!

OpenStudy (jhannybean):

Is your question: \(7 \cdot 3^3 x-2=21\)? or is it \(7\cdot 3^{3x} -2=21\)?

OpenStudy (anonymous):

\[7 .3^{3x-2} = 21\]

OpenStudy (anonymous):

7 multiplied by.....

OpenStudy (jhannybean):

Oh alright

OpenStudy (jhannybean):

Divide both sides by 7 first.

OpenStudy (jhannybean):

What does your equation become?

OpenStudy (jhannybean):

Once you do that, you'll see that you have the same base, and therefore can evaluate the powers.